OpenStudy (anonymous):

As cars passed a checkpoint, the following speeds were clocked and recorded. Speed (mph): 55 62 61 54 68 72 59 61 70 Find the mean and the range of the data set.

OpenStudy (jack1):

range = 72 - 54...? so range = 18...?

OpenStudy (jack1):

mean = sum of all results / number of results ...? =562 / 9...? = 62.4... ish???

OpenStudy (anonymous):

Thnx can you help me with my next question?

OpenStudy (jack1):

sure, what is it?